The weight of corn chips dispensed into a 24-ounce bag by the dispensing machine has been identified as possessing a normal distribution with a mean of 24.5 ounces and a standard deviation of 0.2 ounce. What proportion of the 24-ounce bags contain more than the advertised 24 ounces of chips?
Answer #1

Solution :

Given that ,

mean = = 24.5

standard deviation = = 0.2

P(x > 24) = 1 - P(x < 24)

= 1 - P((x - ) / < (24 - 24.5) / 0.2)

= 1 - P(z < -2.5)

= 1 - 0.0062

= 0.9938

Proportion = 0.9938